OpenStudy (anonymous):

Please help! Half of a number increased by two and two thirds is equal to the number decreased by a quarter. What is the number?

hartnn (hartnn):

number =x Half of a number = x/2 two and two thirds =2 2/3 = 7/3 number decreased by a quarter = x-1/4 now can u form the equation ??

OpenStudy (anonymous):

So would it be x/2+ 2 2/3= x-1/4?

hartnn (hartnn):

yes, can u solve that ?

OpenStudy (anonymous):

yes right bro.........

OpenStudy (anonymous):

ok I got it thank you!

hartnn (hartnn):

welcome ^_^